find vector xy that goes from x(8,2,-9) and y (-12,-1,10)

Answer by tag22man(4)

Given XY = XO + OY (O being the origin given vectors and points RELATIVE to the origin)
Then XY = -OX + OY
Thus XY = OY - OX , where OY is simply y (-12,-1,10) and OX is x(8,2,-9)
Therefore using vector subtraction
XY = (-20,3,-19)